It is assumed that consumption C depends on income Y and that this relationship takes the form of the linear function C= a + bY. When Y is N100, C' is observed to be N305. When Y is N150, C is observed to be N455. What are the values of a and b in this function?
1
Expert's answer
2021-10-24T18:14:57-0400

"305 = a + b100 \\

455 = a +b150 \\

a = 305-100b \\

455 = 305 -100b+150b \\

455-305=150b-100b \\

150 = 50b \\

b= \frac{150}{50}=3 \\

305 = a + 3 \times 100 \\

305 = a +300 \\

a = 305-300 = 5"

So, a= 5 and b=3